Volker is an employee of Equus Corporation. He also has his own business working as a ferrier. For 2019, Volker's wages from Equus were $148,000. His self-employment income was $36,500.

If required, round any computations to two decimal places and use in subsequent computations. Round final answers to the nearest dollar.

If an amount is zero, enter "0".

a. Compute the total payroll, self-employment, and additional Medicare taxes for Volker for 2019.

b. Assume the same facts except that Volker's 2019 self-employment activities resulted in a loss of $15,000. Compute the payroll, self-employment, and additional Medicare taxes for Volker for 2019.

c. Compute Volker's NIIT for 2019 using wages of $148,000, self-employment income of $36,500, and NII of $10,000.
Answer #1

a.

  • Payroll, which is the combination of Social Security tax rate (6.2%) and the Medicare tax rate (1.45%). the maximum wage base is 132900, which is less than Volker´s.

132900*7.65=10,166.85 this is the payroll tax for 2019.

  • For Self-employment, only 92.35% of the net earnings from are subject to self-employment tax.

36500*92.35%=33,707

Self employment taxes= 33707*15.35%=5157.28

  • The additional Medicare tax doesn´t apply to Volker because his net earnings are less than 200,000OMB No. 1545-0074 SCHEDULE SE (Form 1040 or 1040-SR) Self-Employment Tax 2019 Attachment Sequence No. 17 Department of the Tr